OpenStudy (anonymous):

Given that P = (5, 9) and Q = (13, 12), find the component form and magnitude of PQ. helppp please

OpenStudy (raden):

first, PQ = vector q - vector p = (13,12) - (5,9) = (8,3) the magnitude of PQ = absolute(x^2+y^2) just put x=8 and y=3

OpenStudy (anonymous):

can i ask anither question. Two forces with magnitudes of 150 and 75 pounds act on an object at angles of 30° and 150° respectively. Find the direction and magnitude of these forces. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

do i use the law of cosines?

OpenStudy (anonymous):

i think you can use the law of cosine

OpenStudy (anonymous):

to find c\[^{2}\] and what angle would i use?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ahm wait i need to check it first